# Joint moments of characteristic polynomials from the orthogonal and unitary symplectic groups

### v1

@@ 1.1 Background and motivation
In this paper we address the problem of calculating joint moments of the characteristic polynomials, along with their higher-order derivatives, of random matrices drawn from the unitary symplectic $\mathrm{USp}(2N)$ and even orthogonal $\mathrm{SO}(2N)$ classical compact groups, establishing a connection with Painlevé equations. The general question can be phrased as follows.
@@
Expressions for the coefficients $(\mathsf{t}_{a}^{(k)})_{k=1}^{\infty}$ in terms of multiple integrals involving Brownian motions have been given in [51] and it would be desirable to understand if there are connections to the moment formulae we provide below.
@@ 1.2 Asymptotics of the joint moments and the Bessel point process
Let a>1a>-1. We define the Bessel point process on $\mathbb{R}_{+}$ with parameter $a$ to be the determinantal point process [15, 45, 34] with correlation kernel $\mathsf{K}_{a}^{\mathrm{Bes}}$, see [33, 34],
$\mathsf{K}_{a}^{\mathrm{Bes}}(x,y)=2\frac{\sqrt{x}J_{a+1}(\sqrt{x})J_{a}(\sqrt{y})-\sqrt{y}J_{a+1}(\sqrt{y})J_{a}(\sqrt{x})}{x-y},$
where $J_{a}$ denotes the Bessel function with parameter $a$. It almost surely consists of a strictly increasing sequence [^1] We note that the superscript $l$ in $\mathfrak{B}_{a}^{l}$ does not correspond to raising to the $l$-th power but rather denotes the $l$-th ordered point in the point process. $0<\mathfrak{B}_{a}^{1}<\mathfrak{B}_{a}^{2}<\mathfrak{B}_{a}^{3}<\cdots$.
@@
$\lim_{N\to\infty}\frac{\mathfrak{J}^{\mathrm{USp}}_{N}(h_{0},h_{1},\ldots,h_{m})}{N^{\frac{s(s+1)}{2}+\sum_{k}kh_{k}}}=\frac{2^{\frac{s^{2}}{2}}G(1+s)\sqrt{\Gamma(1+s)}}{\sqrt{G(1+2s)\Gamma(1+2s)}}\mathbb{E}\left[\prod_{k=1}^{m}\mathfrak{R}_{k}\bigg{(}s+\frac{1}{2}\bigg{)}^{h_{k}}\right]<\infty$
@@
$\lim_{N\to\infty}\frac{\mathfrak{J}_{N}^{\mathrm{SO}}(h_{0},h_{1},\ldots,h_{m})}{N^{\frac{s(s-1)}{2}+\sum_{k}kh_{k}}}=\frac{2^{\frac{s^{2}}{2}}G(1+s)\sqrt{\Gamma(1+2s)}}{\sqrt{G(1+2s)\Gamma(1+s)}}\mathbb{E}\left[\prod_{k=1}^{m}\mathfrak{R}_{k}\bigg{(}s-\frac{1}{2}\bigg{)}^{h_{k}}\right]<\infty.$ (2)
@@
$\mathbb{E}\left[\prod_{k=1}^{m}\left(\mathfrak{e}_{k}(a)\right)^{h_{k}}\right]=\frac{(1!)^{2h_{1}}\cdots(m!)^{2h_{m}}}{Q!}\sum_{j=m}^{Q}(-1)^{Q+j}{Q\choose j}\mathbb{E}_{j}^{(a)}\left[\prod_{k=1}^{m}\left(\mathrm{e}_{k}\left(\frac{1}{\mathsf{x}_{1}^{(j)}},\frac{1}{\mathsf{x}_{2}^{(j)}},\ldots,\frac{1}{\mathsf{x}_{j}^{(j)}}\right)\right)^{h_{k}}\right].$
@@
Following these heuristics and using the results of this paper, we can give conjectures regarding the asymptotic mean value distribution of derivatives of arbitrary order of families of $L$-functions at the central point $s=1/2$. In the following, we present an example conjecture in the even orthogonal case. Theorem 1.3, together with the recipe for formulating conjectures about the moments of any primitive family of $L$-functions, enables similar conjectures to be made for other families of $L$-functions. We refer the reader to [23, Sections 1.3] and [22, Section 4.3] for some examples without derivatives. Consider the mean values of
@@
at the critical point $s=\frac{1}{2}$, averaged over $f\in H_{m}^{*}(N)$. Here, $H_{m}^{*}(N)$ denotes the set of primitive Hecke eigenforms of weight $m$ relative to the subgroup $\Gamma_{0}(N)$, and $\lambda_{f}(n)$ are the Fourier coefficients of the eigenform. For simplicity, we restrict our attention to the case $m=2$ and $N=q$, where $q$ is a prime number. The low-lying zeros of this family near the critical point exhibit orthogonal symmetry, and the functional equation for the $L$-function can be written as,
$\Lambda_{f}(s)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Bigg{(}\sqrt{\frac{2\pi}{q}}\Bigg{)}^{s}\Gamma\left(s+\frac{1}{2}\right)L_{f}(s)=\varepsilon_{f}\Lambda_{f}(1-s),$
@@
Let $k,m\geq 0$ be non-negative integers. Then, as $q\to\infty$,
$\sum_{f~even}\raisebox{3.01385pt}{\hskip-5.11108pt${}^{h}$}\left(L^{(k)}_{f}\Big{(}\frac{1}{2}\Big{)}\right)^{m}\sim a_{s}\big{(}\log(q^{1/2})\Big{)}^{{}^{s(s-1)/2+mk}}g(k;m),$
@@
$g(k;m)=\frac{2^{\frac{m^{2}}{2}-1}G(1+m)\sqrt{\Gamma(1+2m)}}{\sqrt{G(1+2m)\Gamma(1+m)}}\sum_{l_{1},\dots,l_{m}=0}^{k}\left(\prod_{j=1}^{m}\binom{k}{l_{j}}\right)(-2)^{mk-\sum_{j=1}^{m}l_{j}}\mathbb{E}\left[\prod_{j=1}^{m}\mathfrak{R}_{l_{j}}\bigg{(}m-\frac{1}{2}\bigg{)}\right],$
@@ 1.3.1 Connections between higher joint moments and Painlevé equations
where $\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})$ are polynomials in $t_{1}$ whose coefficients are polynomials of $a$, with degree at most $\sum_{\ell=2}^{k}(\ell-1)n_{\ell}$. Moreover, the degree of $t_{1}^{m-1}\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})$ in $t_{1}$ is at most $\sum_{\ell=2}^{k}(\ell-1)n_{\ell}-1$.
@@
$\lim_{N\rightarrow\infty}\frac{\int_{\mathrm{USp}(2N)}\left|\varphi_{\mathbf{A}}^{(4)}(0)\right|^{2}\left|\varphi_{\mathbf{A}}(0)\right|^{s-2}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})}{16N^{\frac{s(s+1)}{2}+8}\int_{\mathrm{USp}(2N)}\left(\varphi_{\mathbf{A}}(0)\right)^{s}\mathrm{d}\mu_{\mathrm{Haar}}({\mathbf{A}})}$
$=\,{\frac{4\,{s}^{6}+56\,{s}^{5}+307\,{s}^{4}+806\,{s}^{3}+967\,{s}^{2}+380\,s+36}{\left(2\,s+1\right)^{2}\left(2\,s+7\right)\left(2\,s+5\right)\left(2\,s+3\right)\left(2\,s-1\right)}}.$
@@
$\lim_{N\rightarrow\infty}\frac{\int_{\mathrm{SO}(2N)}\left|\varphi_{\mathbf{A}}^{(4)}(0)\right|^{2}\left|\varphi_{\mathbf{A}}(0)\right|^{s-2}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})}{16N^{\frac{s(s-1)}{2}+8}\int_{\mathrm{SO}(2N)}\left(\varphi_{\mathbf{A}}(0)\right)^{s}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})}$
$=\,{\frac{4\,{s}^{6}+32\,{s}^{5}+87\,{s}^{4}+58\,{s}^{3}-109\,{s}^{2}-108\,s+72}{\left(2\,s-1\right)^{2}\left(2\,s+5\right)\left(2\,s+3\right)\left(2\,s+1\right)\left(2\,s-3\right)}}.$
@@
For reasons that will become clear in Section 3, joint moments over $\mathrm{SO}(2N)$ and $\mathrm{USp}(2N)$ admit representations in terms of joint moments of the random variables $\{\mathfrak{p}_{m,N}^{(a,b)}\}_{m=1}^{\infty}$. Moreover, it is known that (see [40, 19]) the function
@@
The convergence asserted in Theorem 1.3 for $\mathrm{SO}(2N)$ holds for any $h_{0}\in[0,\infty)$, provided $h_{1},\ldots,h_{m}\in\mathbb{N}.$
@@ 1.3.2 A question of Altuğ, Bettin, Petrow, Rishikesh, and Whitehead
$\det_{1\leq i,j\leq s}\Big{(}f_{2i-j+2l_{i}}(t)\Big{)}$ (11)
for certain non-negative integers $l_{i}$. The approaches used in [1] and [4] are based on the shifted moment formulae of Conrey, Farmer, Keating, Rubinstein, and Snaith in [22], as well as techniques similar to those in [25] by Conrey, Rubinstein, and Snaith. As explained in [1], the authors were motivated by the result in [25], which expressed the leading coefficients of $M_{s}(G(N),1)$ in terms of the derivatives (with respect to $t$ at zero) of
$\det_{1\leq i,j\leq s}\Big{(}I_{i+j-1}(2\sqrt{t})\Big{)}$ (12)
in the case where $G(N)=\mathrm{U}(N)$, where $I_{n}(x)$ is the modified Bessel function of the first kind. In [49], the third and fourth-named authors of this paper proved that, for any $k\geq 2$, $M_{s}(\mathrm{U}(N),k)$ can also be expressed in terms of the derivatives of (12) evaluated at $t=0$. The Hankel determinant (12) was shown by Forrester and Witte [38] to be an Okamoto $\tau$-function associated with a $\sigma$-Painlevé III. In light of all these developments, it becomes natural to consider the following question that was initially posed in [1].
@@
As a direct consequence of our results we provide a positive answer to this question by confirming that a suitable choice of differential equation is the $\sigma$-Painlevé III equation itself, as given in (5). More specifically, our results confirm that in the case considered herein, $\mathbb{E}\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$ plays a role analogous to the role of (12) in the unitary case by leveraging the fact that the function
satisfies the $\sigma$-Painlevé III equation (5), see [7]. In what follows, we give a precise statement: for any integer $k\geq 1$ and for $G(N)=\mathrm{USp}(2N)$ or $\mathrm{SO}(2N)$, the leading coefficients of the moments $M_{s}(G(N),k)$ can be expressed in terms of the derivatives (with respect to $t$ at zero) of $\mathbb{E}\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$. This also holds for $G(N)=\mathrm{O}^{-}(2N)$ by a similar argument. For the sake of brevity, we omit the details.
@@
$\Bigg{(}(-1)^{m}\frac{1}{m!}\frac{\mathrm{d}^{m}}{\mathrm{d}t^{m}}\mathbb{E}\Big{[}\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\Big{]}+(-1)^{m}\sum_{\begin{subarray}{c}j_{-1}+j_{0}+j_{1}+2j_{2}+\cdots+mj_{m}=m\\ j_{2}+j_{3}+\cdots+j_{m}\neq 0\end{subarray}}\frac{1}{\prod_{i=2}^{m}(-i)^{j_{i}}\prod_{i=-1}^{m}j_{i}!}$
@@
with $\mathcal{P}_{n}^{(a)}(j_{2},\ldots,j_{m};t)$ given as in Theorem 1.6 and $\mathbb{E}[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}]$ is the $\tau$ function of the $\sigma$-Painlevé III equation (5).
@@
Note that $\phi_{\mathbf{A}}(\mathrm{e}^{-{\textnormal{i}}\theta})=\varphi_{\mathbf{A}}(\theta)$. So it is not hard to check that the leading coefficient of the moment $M_{s}(G(N),k)$ coincides with that of $\int_{G(N)}\left|\varphi_{\mathbf{A}}^{(k)}(0)\right|^{s}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})$ when $G(N)=\mathrm{USp}(2N)$ or $\mathrm{SO}(2N)$. Hence, the above result is an immediate consequence of combining Theorems 1.3 and 1.6. In fact, these results yield a more general version of Corollary 1.12, describing the joint moments of the $k_{i}$-th derivatives for $i=1,\ldots,m$. To help the reader better understand the result, we present it here for the case $m=1$ only. The results for general $m$ are analogous but have more complicated forms. As a consequence of these results, we can recursively compute all $M_{s}(G(N),k_{1},\ldots,k_{m})$ by using an asymptotic analysis of the Taylor expansion of $\mathbb{E}\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$ at $t=0$, due to the properties of solutions to the corresponding Painlevé equation.
@@
We now briefly explain how the approach we use herein that leads to the proof of Corollary 1.12 is different from those used in [1] and [4]. Initially, in exploring connections between the asymptotic behavior of $M_{s}(G(N),k)$ for unitary symplectic and orthogonal groups and the Painlevé equations, we began with the determinant representation $\det_{1\leq i,j\leq s}\Big{(}f_{2i-j+2l_{i}}(t)\Big{)}$ as in (10). See [39] for some investigations in this direction. However, the lack of a Hankel structure in this determinant posed a fundamental obstacle for further analysis, since the Hankel determinant is the basic object in Okamoto’s $\tau$-function theory and the Riemann-Hilbert method. Therefore, instead of working directly with (10) or (11), we turned to a hidden Hankel structure in $M_{s}(G(N),2)$ for finite $N$, whose entries involve the Laplace transform of certain Jacobi-like weights. For general $k$, we expressed the moment in terms of a Hankel determinant shifted by partitions. We then used combinatorial methods to analyze the connections between these shifted Hankel determinants and the original one, which satisfies a $\sigma$-Painlevé V equation. Finally, after appropriate rescaling and normalization (by dividing by a suitable power of $N$), we first establish the existence of the limit using our results from Section 2, and then by taking the limit as $N\to\infty$ in the expression that connects the moments for finite matrix size $N$ with the $\sigma$-Painlevé V equation, we obtain the desired relation between the leading coefficients of $M_{s}(G(N),k)$ and a solution of the $\sigma$-Painlevé III equation. A more detailed sketch of the proof strategy is provided in Section 4.1. To end this subsection, we present a recurrence relation for the Laplace transform $\mathbb{E}\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$ in order to see the distinction between this and the determinant $\mathcal{T}_{s,l}$ from (10).
Proposition 1.13.
Let $a\in\mathbb{R}$ with $a>1$. Then for $t\geq 0$,
The above recurrence relation (with $a=s+\frac{1}{2}$ for symplectic symmetry and $a=s-\frac{1}{2}$ for even orthogonal symmetry) differs from the recurrence relation for $\mathcal{T}_{s,l}$, which is stated in [1, Theorem 4] as follows:
$\mathcal{T}_{s+1,l}(t)\,\mathcal{T}_{s-1,l}(t)=2\Bigg{(}t\frac{\mathrm{d}^{2}\mathcal{T}_{s,l}(t)}{\mathrm{d}t^{2}}\mathcal{T}_{s,l}(t)+\mathcal{T}_{s,l}(t)\frac{\mathrm{d}\mathcal{T}_{s,l}(t)}{\mathrm{d}t}-t\left(\frac{\mathrm{d}\mathcal{T}_{s,l}(t)}{\mathrm{d}t}\right)^{2}\Bigg{)}$
for $s\in\mathbb{N}$ and $l\in\mathbb{Z}$. Both of the above identities are derived from the Lewis Carroll formula. Obtaining higher order relations, involving several products, of this type is an interesting problem. For the detailed proof of Proposition 1.13, we refer to the paragraph following Proposition 4.1.

### v2

@@ 1.1 Background and motivation
In this paper we address the problems of the convergence of joint moments of the characteristic polynomials with positive real exponents, along with their higher-order derivatives, of random matrices drawn from the unitary symplectic $\mathrm{USp}(2N)$ and even orthogonal $\mathrm{SO}(2N)$ classical compact groups after suitable scalings, and of establishing connections with Painlevé equations for both finite matrix size $N$ and in the large-$N$ limit. The general question can be phrased as follows.
@@
Expressions for the coefficients $(\mathsf{t}_{a}^{(k)})_{k=1}^{\infty}$ in terms of multiple integrals involving Brownian motions have been given in [50]. Below Proposition 1.4, we explain a connection between these coefficients and the joint moments considered in the following section.
@@ 1.2 Asymptotics of the joint moments and the Bessel point process
Let a>1a>-1. We define the Bessel point process on $\mathbb{R}_{+}$ with parameter $a$ to be the determinantal point process [15, 44, 32, 33] with correlation kernel $\mathsf{K}_{a}^{\mathrm{Bes}}$,
$\mathsf{K}_{a}^{\mathrm{Bes}}(x,y)=\frac{\sqrt{x}J_{a+1}(\sqrt{x})J_{a}(\sqrt{y})-\sqrt{y}J_{a+1}(\sqrt{y})J_{a}(\sqrt{x})}{2(x-y)},$
where $J_{a}$ denotes the Bessel function of the first kind with parameter $a$. It almost surely consists of a strictly increasing sequence [^1] We note that the superscript $l$ in $\mathfrak{B}_{a}^{l}$ does not correspond to raising to the $l$-th power but rather denotes the $l$-th ordered point in the point process. $0<\mathfrak{B}_{a}^{1}<\mathfrak{B}_{a}^{2}<\mathfrak{B}_{a}^{3}<\cdots$.
@@
$\lim_{N\to\infty}\frac{\mathfrak{J}^{\mathrm{USp}}_{N}(h_{0},h_{1},\ldots,h_{m})}{N^{\frac{s(s+1)}{2}+\sum_{k=1}^{m}kh_{k}}}=\frac{2^{\frac{s^{2}}{2}}G(1+s)\sqrt{\Gamma(1+s)}}{\sqrt{G(1+2s)\Gamma(1+2s)}}\mathbb{E}\left[\prod_{k=1}^{m}\mathfrak{R}_{k}\bigg(s+\frac{1}{2}\bigg)^{h_{k}}\right]<\infty$
@@
$\lim_{N\to\infty}\frac{\mathfrak{J}_{N}^{\mathrm{SO}}(h_{0},h_{1},\ldots,h_{m})}{N^{\frac{s(s-1)}{2}+\sum_{k=1}^{m}kh_{k}}}=\frac{2^{\frac{s^{2}}{2}}G(1+s)\sqrt{\Gamma(1+2s)}}{\sqrt{G(1+2s)\Gamma(1+s)}}\mathbb{E}\left[\prod_{k=1}^{m}\mathfrak{R}_{k}\bigg(s-\frac{1}{2}\bigg)^{h_{k}}\right]<\infty.$ (2)
@@
$\mathbb{E}\left[\prod_{k=1}^{m}\left(\mathfrak{e}_{k}(a)\right)^{h_{k}}\right]=\frac{(1!)^{h_{1}}\cdots(m!)^{h_{m}}}{Q!}\sum_{j=m}^{Q}(-1)^{Q+j}{Q\choose j}\mathbb{E}_{j}^{(a)}\left[\prod_{k=1}^{m}\left(\mathrm{e}_{k}\left(\frac{1}{\mathsf{x}_{1}^{(j)}},\frac{1}{\mathsf{x}_{2}^{(j)}},\ldots,\frac{1}{\mathsf{x}_{j}^{(j)}}\right)\right)^{h_{k}}\right].$
@@
Following these heuristics and using the results of this paper, we can give conjectures regarding the asymptotic mean value distribution of derivatives of arbitrary order of families of $L$-functions at the central point $s=1/2$. In the following, we present an example conjecture in the even orthogonal case. Theorem 1.3, together with the recipe for formulating conjectures about the moments of any primitive family of $L$-functions, enables similar conjectures to be made for other families of $L$-functions. We refer the readers to [22, Sections 1.3] and [21, Section 4.3] for some examples without derivatives. Consider the mean values of
@@
at the critical point $s=\frac{1}{2}$, averaged over $f\in H_{m}^{*}(N)$. Here, $H_{m}^{*}(N)$ denotes the set of primitive Hecke eigenforms of weight $m$ relative to the subgroup $\Gamma_{0}(N)$, and $\lambda_{f}(n)$ are the normalized Fourier coefficients of the eigenform with $\lambda_{f}(1)=1$. For simplicity, we restrict our attention to the case $m=2$ and $N=q$, where $q$ is a prime number. The low-lying zeros of this family near the critical point exhibit orthogonal symmetry, and the functional equation for the $L$-function can be written as,
$\Lambda_{f}(s)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Bigg(\frac{\sqrt{q}}{2\pi}\Bigg)^{s}\Gamma\left(s+\frac{1}{2}\right)L_{f}(s)=\varepsilon_{f}\Lambda_{f}(1-s),$
@@
Let $k,s\geq 0$ be non-negative integers. Then, as $q\to\infty$,
$\sum_{f~even}\raisebox{3.01385pt}{\hskip-5.11108pt${}^{h}$}\left(L^{(k)}_{f}\Big(\frac{1}{2}\Big)\right)^{s}\sim a_{s}\big(\log(q^{1/2})\Big)^{{}^{s(s-1)/2+sk}}g(k;s),$
@@
$g(k;s)=\frac{2^{\frac{s^{2}}{2}-1}G(1+s)\sqrt{\Gamma(1+2s)}}{\sqrt{G(1+2s)\Gamma(1+s)}}\sum_{l_{1},\dots,l_{s}=0}^{k}\left(\prod_{j=1}^{s}\binom{k}{l_{j}}\right)(-2)^{sk-\sum_{j=1}^{s}l_{j}}\mathbb{E}\left[\prod_{j=1}^{s}\mathfrak{R}_{l_{j}}\bigg(s-\frac{1}{2}\bigg)\right],$
@@ 1.3.1 Connections between higher joint moments and Painlevé equations
where $\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})$ are polynomials in $t_{1}$ whose coefficients are polynomials in $a$ of degree at most $\sum_{\ell=2}^{k}(\ell-1)n_{\ell}$. Moreover, the degree of $t_{1}^{m-1}\mathcal{P}_{m}^{(a)}(n_{2},\ldots,n_{k};t_{1})$ in $t_{1}$ is at most $\sum_{\ell=2}^{k}(\ell-1)n_{\ell}-1$.
@@
$\lim_{N\rightarrow\infty}\frac{\int_{\mathrm{USp}(2N)}\left|\varphi_{\mathbf{A}}^{(4)}(0)\right|^{2}\left|\varphi_{\mathbf{A}}(0)\right|^{s-2}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})}{N^{\frac{s(s+1)}{2}+8}\int_{\mathrm{USp}(2N)}\left(\varphi_{\mathbf{A}}(0)\right)^{s}\mathrm{d}\mu_{\mathrm{Haar}}({\mathbf{A}})}$
$=16{\frac{4\,{s}^{6}+56\,{s}^{5}+307\,{s}^{4}+806\,{s}^{3}+967\,{s}^{2}+380\,s+36}{\left(2\,s+1\right)^{2}\left(2\,s+7\right)\left(2\,s+5\right)\left(2\,s+3\right)\left(2\,s-1\right)}}.$
@@
$\lim_{N\rightarrow\infty}\frac{\int_{\mathrm{SO}(2N)}\left|\varphi_{\mathbf{A}}^{(4)}(0)\right|^{2}\left|\varphi_{\mathbf{A}}(0)\right|^{s-2}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})}{N^{\frac{s(s-1)}{2}+8}\int_{\mathrm{SO}(2N)}\left(\varphi_{\mathbf{A}}(0)\right)^{s}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})}$
$=16{\frac{4\,{s}^{6}+32\,{s}^{5}+87\,{s}^{4}+58\,{s}^{3}-109\,{s}^{2}-108\,s+72}{\left(2\,s-1\right)^{2}\left(2\,s+5\right)\left(2\,s+3\right)\left(2\,s+1\right)\left(2\,s-3\right)}}.$
@@
For reasons that will become clear in Section 3, joint moments over $\mathrm{SO}(2N)$ and $\mathrm{USp}(2N)$ admit representations in terms of joint moments of the random variables $\{\mathfrak{p}_{m,N}^{(a,b)}\}_{m=1}^{\infty}$. Moreover, it is known that (see e.g., [36, Section 3]) the function
@@
The convergence asserted in Theorem 1.3 for $\mathrm{SO}(2N)$ holds for any $h_{0}\in[0,\infty)$, provided $h_{1},\ldots,h_{m}\in\mathbb{N}\cup\{0\}.$
@@ 1.3.2 A question of Altuğ, Bettin, Petrow, Rishikesh, and Whitehead
$\det_{1\leq i,j\leq s}\Big(f_{2i-j-1+l_{i}}(t)\Big)$ (11)
for certain non-negative integers $l_{i}$. The approaches used in [1] and [4] are based on the shifted moment formulae of Conrey, Farmer, Keating, Rubinstein, and Snaith in [21], as well as techniques similar to those in [24] by Conrey, Rubinstein, and Snaith. As explained in [1], the authors were motivated by the result in [24], which expressed the leading coefficients of $M_{s}(\mathrm{U}(N),1)$ in terms of the derivatives (with respect to $t$ at zero) of
$\det_{1\leq i,j\leq s}\Big(I_{i+j-1}(2\sqrt{t})\Big),$ (12)
where $I_{n}(x)$ is the modified Bessel function of the first kind. In [48], the third and fourth-named authors of this paper proved that, for any $k\geq 2$, $M_{s}(\mathrm{U}(N),k)$ can also be expressed in terms of the derivatives of (12) evaluated at $t=0$. The Hankel determinant (12) was shown by Forrester and Witte [37] to be an Okamoto $\tau$-function associated with a $\sigma$-Painlevé III. This connection was later established alternatively by Basor et al. [14] using the Riemann–Hilbert method. In light of all these developments, it becomes natural to consider the following question that was initially posed in [1].
@@
As a direct consequence of our results we provide a positive answer to this question by confirming that a suitable choice of differential equation is the $\sigma$-Painlevé III equation itself, as given in (5). More specifically, our results confirm that, in the case considered here, $\mathbb{E}\!\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$ plays a role analogous to that of (12) in the unitary case, for the following two reasons. On the one hand, $\mathbb{E}\!\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$ is the $\tau$-function associated with the $\sigma$-Painlevé III equation(5). On the other hand, in what follows we give a precise statement: for any integer $k\geq 1$ and for $G(N)=\mathrm{USp}(2N)$ or $\mathrm{SO}(2N)$, the leading coefficients of the moments $M_{s}(G(N),k)$ can be expressed in terms of the derivatives at $t=0$ of $\mathbb{E}\!\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right].$ By a similar argument, the same conclusion also holds for $G(N)=\mathrm{O}^{-}(2N)$. For the sake of brevity, we omit the details.
@@
$\Bigg((-1)^{m}\frac{1}{m!}\frac{\mathrm{d}^{m}}{\mathrm{d}t^{m}}\mathbb{E}\Big[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\Big]+(-1)^{m}\sum_{\begin{subarray}{c}j_{-1}+j_{0}+j_{1}+2j_{2}+\cdots+mj_{m}=m\\ j_{2}+j_{3}+\cdots+j_{m}\neq 0\end{subarray}}\frac{(-1)^{j_{-1}}}{\prod_{i=2}^{m}(-i)^{j_{i}}\prod_{i=-1}^{m}j_{i}!}$
@@
with $\mathcal{P}_{n}^{(a)}(j_{2},\ldots,j_{m};t)$ given as in Theorem 1.6 and
satisfies the $\sigma$-Painlevé III equation (5).
@@
The above result is an immediate consequence of combining Theorems 1.3 and 1.6. In fact, note that $\phi_{\mathbf{A}}(\mathrm{e}^{-{\textnormal{i}}\theta})=\varphi_{\mathbf{A}}(\theta)$, so it is not hard to check that the leading coefficient of the moment $M_{s}(G(N),k)$ coincides with that of $\int_{G(N)}\left|\varphi_{\mathbf{A}}^{(k)}(0)\right|^{s}\mathrm{d}\mu_{\mathrm{Haar}}(\mathbf{A})$ when $G(N)=\mathrm{USp}(2N)$ or $\mathrm{SO}(2N)$. Moreover, these results yield a more general version of Corollary 1.12, describing the joint moments of the $k_{i}$-th derivatives for $i=1,\ldots,m$. To help the reader better understand the result, we present it here for the case $m=1$ only. The results for general $m$ are analogous but have more complicated forms. As a consequence of these results, we can recursively compute all $M_{s}(G(N),k_{1},\ldots,k_{m})$ by using an asymptotic analysis of the Taylor expansion of $\mathbb{E}\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$ at $t=0$, due to the properties of solutions to the corresponding Painlevé equation.
We now briefly explain how the approach we use herein that leads to the proof of Corollary 1.12 is different from those used in [1] and [4]. Initially, in exploring connections between the asymptotic behavior of $M_{s}(G(N),k)$ for unitary symplectic and orthogonal groups and the Painlevé equations, we began with the determinant representation $\det_{1\leq i,j\leq s}\Big(f_{2i-j-1+l_{i}}(t)\Big)$ as in (10). See [38] for some investigations in this direction. However, the absence of a Hankel structure or a Toeplitz property in this determinant poses a challenge for further analysis through Okamoto’s $\tau$-function theory or the Riemann-Hilbert method.
@@
Therefore, instead of working directly with (10) or (11), we turned to a hidden Hankel structure in $M_{s}(G(N),2)$ for finite $N$, whose entries involve the Laplace transform of certain Jacobi-like weights. For general $k$, we expressed the moment in terms of a Hankel determinant shifted by partitions. We then used combinatorial methods to analyze the connections between these shifted Hankel determinants and the original one, which satisfies a $\sigma$-Painlevé V equation. Finally, after appropriate rescaling and normalization (by dividing by a suitable power of $N$), we first establish the existence of the limit using our results from Section 2, and then by taking the limit as $N\to\infty$ in the expression that connects the moments for finite matrix size $N$ with the $\sigma$-Painlevé V equation, we obtain the desired relation between the leading coefficients of $M_{s}(G(N),k)$ and a solution of the $\sigma$-Painlevé III equation. A more detailed sketch of the proof strategy is provided in Section 4.1.